OpenStudy (anonymous):

Sandy purchases a new car for $18,000. The car depreciates (loses value) at a rate of 8% each year. A. If you wish to trade in the car in 5 years, what is the car worth? B. How many years will it take the car's value to depreciate below $10,000 C. Write a general formula that could be used to find the car's value after n number of years.

OpenStudy (anonymous):

8% of 18000 = 1440 the car's value decreases by $1440 each year. a/ 1440 * 5 = 7200 18000 - 7200 = 10800 b/ 6 c/ y = -1440x + 18000 x is the time, y is the value